OpenStudy (anonymous):

A baseball manager must choose a 9-man batting order from a team of 25 players. How manu possible batting orders are there?

OpenStudy (anonymous):

it doesn't say so (?)

OpenStudy (dumbcow):

\[=\frac{25!}{9!}\]

OpenStudy (dumbcow):

it a batting order, so yes order matters use permutations 25P9

OpenStudy (dumbcow):

oops i messed up, it should be 16! in denominator

OpenStudy (dumbcow):

\[nPr = \frac{n!}{(n-r)!}\]

OpenStudy (anonymous):

25!/16! gives me 741355000000?

OpenStudy (dumbcow):

yeah its a huge number thats how many possible lineups there are crazy huh

OpenStudy (anonymous):

wowsers :D

OpenStudy (anonymous):

can i ask you another on here or should i start a new thread?

OpenStudy (dumbcow):

no go ahead

OpenStudy (anonymous):

it says: of the first ten questions on a quiz, a student must answer 7. of the second 5 questions a student must answer 3. in how many different ways can this be done?

OpenStudy (anonymous):

so......10!/3! x 5!/2!?

OpenStudy (dumbcow):

yes you have right idea however in this case order doesn't matter. it doesn't matter which 7 of the 10 they answer so use combinations \[nCr = \frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

whoa

OpenStudy (dumbcow):

you should end up with 1200

OpenStudy (anonymous):

workin it